OpenStudy (anonymous):

Please help -3/2x - 1 = 2/5x + 4

OpenStudy (anonymous):

Is that (-3/2)x and (2/5)x or -3 over 2x and 2 over 5x?

OpenStudy (anonymous):

(-3/2)x - 1 = (2/5)x + 4

OpenStudy (anonymous):

You can get rid of the fractions if you multiply through by a factor of 10. If you multiply every term on both sides of the equation, you don't change the equation. So, multiplying (-3/2)x by 10, you get -15x multiplying (2/5)x by 10 you get 4x multiplying -1 by 10 results in -10 and multiplying 4 by 10 results in 40, so you end up with -15x-10=4x+40 Can you take it from there, or do you need more help?

OpenStudy (anonymous):

Of course, I only chose 10 because it worked for this equation. Other equations would need a different factor.

OpenStudy (anonymous):

need more help.

OpenStudy (anonymous):

not understanding how you got 4x from 2/5 x 10

OpenStudy (anonymous):

2/5 x 10 = 20/5 20 divided by 5 is 4

OpenStudy (anonymous):

-3/2 x 10 = -30/2 -30 divided by 2 is -15

OpenStudy (anonymous):

me have no idea

OpenStudy (anonymous):

I have never seen it work that way by multiplying everything by 10

OpenStudy (anonymous):

so am I going to add -15x and 4x or subtract

OpenStudy (gw2011):

-(3/2)x-1=(2/5)x+4 Combining like terms together you get: -(3/2)x-(2/5)x=4+1 -(3/2)x-(2/5)x=5 Multiply by 10 in order to eliminate the denominators 2 & 5: -15x-4x=50 -19x=50 Divide by -19 to find x: x=-50/19

OpenStudy (anonymous):

ok. Thank you guys!

OpenStudy (gw2011):

You're welcome

OpenStudy (anonymous):

(-3/2)x-1=(2/5)x+4 -1.5x-1=0.4x+4 -1.5x-0.4x=4+1 (-1.5-0.4)x=5 -1.9x=5 x=5/1.9 x=-2.63